# HTD-Mamba: Efficient Hyperspectral Target Detection with Pyramid State Space Model

### v1

@@ Abstract
Hyperspectral target detection (HTD) identifies objects of interest from complex backgrounds at the pixel level, playing a vital role in Earth observation. However, HTD faces challenges due to limited prior knowledge and spectral variations, leading to underfitting models and unreliable performance. To address these challenges, this paper proposes an efficient self-supervised HTD method with a pyramid state space model (SSM), named HTD-Mamba, which employs spectrally contrastive learning to distinguish between target and background based on the similarity measurement of intrinsic features. Specifically, to obtain sufficient training samples and leverage spatial contextual information, we propose a spatial-encoded spectral augmentation technique that encodes all surrounding pixels within a patch into a transformed view of the central pixel. Additionally, to explore global band correlations, we divide pixels into continuous group-wise spectral embeddings and introduce Mamba to HTD for the first time to model long-range dependencies of the spectral sequence with linear complexity. Furthermore, to alleviate spectral variation and enhance robust representation, we propose a pyramid SSM as a backbone to capture and fuse multiresolution spectral-wise intrinsic features. Extensive experiments conducted on four public datasets demonstrate that the proposed method outperforms state-of-the-art methods in both quantitative and qualitative evaluations. Code is available at https://github.com/shendb2022/HTD-Mamba.
@@ I Introduction
Hyperspectral target detection (HTD), a critical downstream task, aims to identify and locate objects of interest based on spectral and spatial discrepancies. Unlike typical “box” detection methods used in RGB images, which delineate objects with various rectangular or rotated boxes, HTD distinguishes between background and target information at the pixel or sub-pixel level. This fine-grained detection and joint spectral-spatial discrimination make HTD highly relevant in current Earth observation and quantitative remote sensing tasks[6, 7]. However, it faces two longstanding challenges:
@@
Over the past decades, various approaches have been developed to address these challenges. Early statistical-based methods [8, 9] assumed certain background distributions to determine target pixels. However, their performance may be limited when applied to real-world datasets due to inconsistencies with prior assumptions. Subsequent representation-based approaches [10, 11] transformed the detection task into an optimization problem based on the linear mixing model (LMM) and various regularizations. Despite their physical interpretability, these models may encounter difficulties due to inaccurate regularization constraints and nonlinear representations.
@@
In contrast, deep learning, with its powerful feature extraction and parallel computation capabilities, has demonstrated effectiveness and superiority in HTD. Due to limited prior knowledge and class imbalance between targets and backgrounds, two categories of deep detectors have emerged: supervised classification methods [12, 13] and unsupervised detection methods [14, 15]. The former expands the training samples via data augmentation and employs a similarity-dissimilarity binary classification model for detection. The latter extracts discriminative intrinsic features through a carefully designed self-supervised model and then separates the target from the background based on these features. Although these methods have achieved considerable performance, several problems still exist. Firstly, it is challenging for supervised classification methods to train an accurate binary classifier due to the distribution mismatch between the expanded samples and the original samples. Secondly, it is difficult for unsupervised detection methods to design a discrimination loss that ensures effective background-target separation. Thirdly, it is demanding for both categories of methods to develop an efficient backbone to obtain robust intrinsic feature representations for HTD.
@@
Current backbones for HTD are mainly based on convolutional neural networks (CNNs) and Transformers. CNN-based methods [12, 16] excel at extracting local features but suffer from a limited receptive field, making it difficult to capture non-local correlations. In contrast, Transformer-based methods [17, 18, 19] can capture global contextual dependencies based on the multihead self-attention mechanism. However, the computational complexity of Transformers is quadratic to the number of embedding tokens, leading to substantial computational overload and memory requirements as the sequence grows. Recently, the state space model (SSM) [20] has gained attention for modeling long sequences with near-linear complexity. Notably, Mamba [21] introduces a selective state space (S6) that makes SSM parameters input-aware and allows selective propagation or forgetting of information based on the current token. Additionally, a hardware parallel scan algorithm in recurrent mode accelerates the training and inference process. Therefore, considering the high-dimensional sequence nature of hyperspectral data, applying Mamba as a backbone for HTD shows promise in efficiently capturing long-range dependencies.
@@
Inspired by the above insights, we propose an efficient HTD method with a pyramid SSM (HTD-Mamba). This method employs spectrally contrastive learning to recognize pixel-wise instances by maximizing the feature similarity between each pixel and its transformed view. Thus, detection can be performed by feeding pairs consisting of each detected pixel and the prior target spectrum into the well-trained model. Specifically, an efficient data augmentation technique is proposed to obtain sufficient training samples and leverage spatial information. This technique encodes all surrounding pixels within a patch into a new spectral view of the central pixel, using spectral similarity as weights. Additionally, to explore global band correlations, the input pixels are dynamically divided into continuous group-wise spectral embeddings using a 1D CNN. The efficient architecture, Mamba, is first introduced to HTD to model long-range dependencies of the spectral sequence with linear complexity. Furthermore, to enhance robust feature representations, a pyramid SSM is developed as a backbone to extract multiresolution spectral-wise intrinsic features by capturing the global correlation of spectral sequences at different resolutions and fusing features from different levels. Therefore, HTD-Mamba effectively addresses the challenges of limited prior knowledge by constructing sufficient view pairs and mitigates spectral variation by extracting multiresolution discriminative intrinsic features using the pyramid SSM. Experimental results across multiple datasets demonstrate the effectiveness and superiority of HTD-Mamba compared to existing state-of-the-art methods.
@@
  2. We propose an efficient data augmentation technique to create sufficient view pairs and leverage spatial contextual information, which encodes all surrounding pixels within a patch into a transformed view of the central pixel.
@@ II-A1 SSM
The SSM is known as a linear time-invariant system that maps a 1D sequence $x(t)\in\mathbb{R}$ to its response $y(t)\in\mathbb{R}$. Mathematically, this process can be formulated using the following linear ordinary differential equations (ODEs):
@@
where $h(t)\in\mathbb{R}^{L}$ denotes the hidden state, ${\bf A}\in\mathbb{R}^{L\times L}$ is the state transition matrix derived from the HiPPO operator to memory the history and capture long-range dependencies, and ${\bf B}\in\mathbb{R}^{L\times 1}$ and ${\bf C}\in\mathbb{R}^{1\times L}$ are the projection matrices to control the input and output of the system, respectively.
@@ II-A2 Discretization
To integrate the continuous SSM into deep learning algorithms, the discretization process is necessay. Specifically, ODEs (1) can be discretized using a zero-order hold rule as follows:
$\begin{split}h_{k}&=\bar{{\bf A}}h_{k-1}+\bar{{\bf B}}x_{k},\\ y_{k}&={\bf C}h_{k}\end{split}$ (2)
@@
$\begin{split}\bar{{\bf A}}&={\rm exp}({\bf\Delta}{\bf A}),\\ \bar{{\bf B}}&=({\bf\Delta}{\bf A})^{-1}({\rm exp}({\bf\Delta}{\bf A}-{\bf I})\cdot{\bf\Delta}{\bf B}\end{split}$ (3)
@@ II-A3 Selective Scan Mechanism
Due to the time-invariant property, i.e., $\bf B$, $\bf C$, and ${\bf\Delta}$ are independent of the input sequence, the output can be calculated either by recurrance or by global convolution. However, this restricts content awareness. To address this, the selective scan mechanism is incorporated, where the projection and time scale matrices are derived from the input, ensuring awareness of the contextual information:
$\begin{split}S_{\bf B}(x)&={\rm Linear}_{N}(x),\\ S_{\bf C}(x)&={\rm Linear}_{N}(x),\\ S_{\bf\Delta}(x)&={\rm Broadcast}_{D}({\rm Linear}_{1}(x))\end{split}$ (4)
where ${\rm Linear}_{d}(x)$ is the parameterized projection to dimension $d$, ${\rm Broadcast}_{d}$ is the broadcast operation to dimension $d$, and $N$ and $D$ are the embedding size and the hidden state size, respectively. Thus, the system becomes time-varying, and only recurrence can be adopted. Additionally, with the incorporation of parallel scan and integration with the gated multilayer perceptron (MLP), SSM evolves into the popular Mamba, which excels at capturing long-range dependencies with near-linear time complexity and hardware acceleration.
Figure 1: The overall architecture of the proposed HTD-Mamba.
Figure 2: Flowchart of pyramid SSM.
@@ II-B Overall Architecture
To address the issues of limited training samples and spectral variation, we propose a self-supervised spectrally contrastive learning method based on Mamba with pyramid SSM. It aims to recognize pixel-level instances by maximizing the similarity and consistency between each pixel and its transformed view. Thus, targets of interest can be detected by pairing them with the prior target spectrum. Fig. 1 illustrates the overall architecture of the proposed method.
The training stage encompasses the spatial-encoded spectral augmentation (SESA), the multiresolution spectral feature extraction (MSFE), and the spectral contrastive head (SCH). The SESA block first creates a new spectral view of the center pixel by encoding the contextual information within a patch, enabling sufficient view pairs for contrastive learning. Then, the MSFE block divides the pixel into continuous group-wise spectral embeddings and treats them as a spectral sequence. To obtain discriminative intrinsic feature representation, a pyramid SSM is proposed to capture multiresolution long-range dependencies of sequences. Finally, the paired extracted features pass through the SCH block to calculate the similarity, and the contrastive loss is applied to ensure model discrimination by pulling the positive pairs close while pushing the negative ones far away.
In the testing stage, each detected pixel is paired with the prior target spectrum and then passes through the MSFE and SCH blocks to obtain the discriminative intrinsic feature vectors. The detection result is obtained by calculating the cosine similarity as
$\mu_{i}=\frac{f({\bf X}_{i})^{T}f({\bf d})}{\|f({\bf X}_{i})\|_{2}\cdot\|f({\bf d})\|_{2}}$ (5)
where ${\bf X}\in\mathbb{R}^{n\times B}$ denotes the image matrix, ${\bf d}\in\mathbb{R}^{B}$ denotes the target spectrum, and $f$ is the backbone consisting of MSFE and SCH. Finally, a nonlinear background suppression function (NBS) is applied to suppress the background:
@@ II-C Spatial-Encoded Spectral Augmentation
Given a random patch ${\bf X}_{p}\in\mathbb{R}^{p^{2}\times B}$, where $p^{2}$ and $B$ denote the number of pixels and bands, respectively, the center pixel of ${\bf X}_{p}$ is ${\bf y}\in\mathbb{R}^{B}$. By encoding all the pixels within the patch using a similarity weight matrix, the new spectral view of ${\bf y}$ can be expressed as
${\bf x}={\bf X}_{p}^{T}{\bf W}$ (7)
where ${\bf W}\in\mathbb{R}^{p^{2}\times 1}$ is the weight matrix measuring the similarity between the center pixel and each pixel within the patch. The weights are defined as
${\bf W}_{i}=\frac{{\rm exp}(C_{i})}{\sum\limits_{j=1}^{p^{2}}{\rm exp}(C_{j})}.$ (8)
The spectral similarity $C_{i}$ is calculated by
$C_{i}=\frac{{\bf y}^{T}{\bf X}_{p}^{i}}{\|{\bf y}\|_{2}\cdot\|{\bf X}_{p}^{i}\|_{2}}$ (9)
where $\|\cdot\|_{2}$ denotes the L2 norm, and ${\bf X}_{p}^{i}$ is the $i$-th pixel of ${\bf X}_{p}$. After obtaining the new spectral view of the center pixel, pairs $({\bf x},{\bf y})$ from the same location comprise the positive samples while pairs from different locations comprise the negative samples. As a result, sufficient training samples can be obtained for contrastive learning.
@@ II-D Multiresolution Spectral Feature Extraction
Specifically, we use 1D convolution with stride to extract local features and divide the spectrum into group representations:
@@
where ${\rm Conv}$ denotes the 1D convolution, $m$ denotes the kernel size, $s=\lceil m/4\rceil$ denotes the stride, $N$ denotes the number of kernels, and $\sigma_{1}$ is the Leaky ReLU function for nonlinear representation. After this process, paired spectral samples $({\bf x},{\bf y})$ can be converted to group-wise embeddings $({\bf S}_{x}\in\mathbb{R}^{L\times N},{\bf S}_{y}\in\mathbb{R}^{L\times N})$, where $L$ denotes the number of spectral groups, satisfying $L=\lfloor(B-m)/s\rfloor+1$. Thus, numerous spectral tokens can be obtained to extract discriminative intrinsic features.
@@
The flowchart of the proposed pyramid SSM is shown in Fig. 2. It comprises bottom-up downsampling, global spectral feature extraction, top-down upsampling and feature fusion, and the gated block. In each deep layer, the input sequence first undergoes a normalization operation to reduce the internal covariate shift and stabilize the training process. In this work, RMSNorm[22] is applied to improve the computational efficiency by eliminating the mean calculation required in Layer Norm:
@@
where ${\rm RMS}$ denotes the root mean square function, ${\bf S}_{i-1}^{j}$ denotes the $j$th channel of the $i$th layer input sequence, and ${\bf g}$ denotes the learnable weight. Then, the normalized sequence is projected into two representations for feature extraction and gated operation, using the fully connected (FC) layer, respectively:
@@
In the bottom-up downsampling process, the temporal sequence is gradually downsampled to obtain multiresolution representations, which facilitates comprehensive analysis. Specifically, the adjacent hierarchical sampling is performed using convolution with stride:
@@
where ${\bf Z}_{1}^{0}={\bf Z}_{1}$, $k\in\{1,2,3\}$ denotes the $k$th pyramid level. The kernel size is fixed at 3, the stride is 2 to perform 2$\times$ downsampling, and the token size is doubled to maintain information capacity.
@@
where ${\rm DWConv}$ denotes the 1D depth-wise convolution, the kernel size is 3, the number of groups is $2^{k+1}N$, $\sigma_{2}$ denotes the SiLU function, and padding is used in the convolution to keep the size fixed.
In the top-down upsampling and feature fusion processes, low-level features are recovered from the adjacent high level and then fused with the extracted global features from the same level. This process can be expressed as
@@
where ${\hat{\bf Z}_{1}^{3}}={\bar{\bf Z}_{1}^{3}}$, ${\rm DConv}$ denotes the transposed 1D convolution with a kernel size of 3 and a stride of 2 to perform 2×2$\times$2 × upsampling. The token size is halved to maintain information capacity. The FC layer is used to perform flexible linear transformation in preparation for fusion.
@@
6:${\bar{\bf A}}$: $(P,L,N,D)$leftarrow ${\bf\Delta}_{o}\otimes{\rm Parameter}^{\bf A}$
7:${\bar{\bf B}}$: $(P,L,N,D)$leftarrow ${\bf\Delta}_{o}\otimes{\bf B}$
@@
13:end for
14:/* Upsampling and feature fusion */
15:${\hat{\bf Z}_{1}^{3}}={\bar{\bf Z}_{1}^{3}}$
16:for $k=3$ to 1 do
17:     ${\hat{\bf Z}_{1}^{k-1}}$: $(P,L_{k-1},2^{k}N)$leftarrow ${\rm DConv}({\hat{\bf Z}_{1}^{k}}))$
18:     ${\hat{\bf Z}_{1}^{k-1}}$: $(P,L_{k-1},2^{k}N)$leftarrow ${\hat{\bf Z}_{1}^{k-1}}+{\rm Linear}_{2^{k}N}({\bar{\bf Z}_{1}^{k-1}})$
19:end for
20:/* Gated block */
21:${\bf Z}$: $(P,L,2N)$leftarrow ${\hat{\bf Z}_{1}^{0}}\odot\sigma_{2}({\bf Z}_{2})$
22:/* Residual connection */
23:${\bf S}_{i}$: $(P,L,N)$leftarrow ${\bf S}_{i-1}+{\rm Linear}_{N}({\bf Z})$
24:${\rm Return}:{\bf S}_{i}$
@@
An overview of the proposed pyramid SSM is summarized in Algorithms 1 and 2.

### v2

@@ Abstract
Hyperspectral target detection (HTD) identifies objects of interest from complex backgrounds at the pixel level, playing a vital role in Earth observation. However, HTD faces challenges due to limited prior knowledge and spectral variation, leading to underfitting models and unreliable performance. To address these challenges, this paper proposes an efficient self-supervised HTD method with a pyramid state space model (SSM), named HTD-Mamba, which employs spectrally contrastive learning to distinguish between target and background based on the similarity measurement of intrinsic features. Specifically, to obtain sufficient training samples and leverage spatial contextual information, we propose a spatial-encoded spectral augmentation technique that encodes all surrounding pixels within a patch into a transformed view of the center pixel. Additionally, to explore global band correlations, we divide pixels into continuous group-wise spectral embeddings and introduce Mamba to HTD for the first time to model long-range dependencies of the spectral sequence with linear complexity. Furthermore, to alleviate spectral variation and enhance robust representation, we propose a pyramid SSM as a backbone to capture and fuse multiresolution spectral-wise intrinsic features. Extensive experiments conducted on four public datasets demonstrate that the proposed method outperforms state-of-the-art methods in both quantitative and qualitative evaluations. Code is available at https://github.com/shendb2022/HTD-Mamba.
@@ I Introduction
Hyperspectral target detection (HTD), a critical downstream task, aims to identify and locate objects of interest based on spectral and spatial discrepancies. Unlike typical “box” detection methods used in RGB images, which delineate objects with various rectangular or rotated boxes, HTD distinguishes between background and target information at the pixel or sub-pixel level. This fine-grained detection and joint spectral-spatial discrimination make HTD highly relevant in current Earth observation and quantitative remote sensing tasks[6, 7, 8]. However, it faces two longstanding challenges:
@@
Over the past decades, various approaches have been developed to address these challenges. Early statistical-based methods [9, 10, 11, 12, 13] assumed certain background distributions to determine target pixels. However, their performance may be limited when applied to real-world datasets due to inconsistencies with prior assumptions. Subsequent representation-based approaches [14, 15, 16, 17, 18] transformed the detection task into an optimization problem based on the linear mixing model and various regularizations. Despite their physical interpretability, these models may encounter difficulties due to inaccurate regularization constraints and weak nonlinear representations.
@@
In contrast, deep learning, with its powerful feature extraction and parallel computation capabilities, has demonstrated effectiveness and superiority in HTD. Due to limited prior knowledge and class imbalance between targets and backgrounds, two categories of deep detectors have emerged: supervised classification methods [19, 20, 21, 22, 23] and unsupervised detection methods [24, 25, 26, 27, 28]. The former expands the training samples via data augmentation and employs a similarity-dissimilarity binary classification model for detection. The latter extracts discriminative intrinsic features through a carefully designed self-supervised model and then separates the target from the background based on these features. Although these methods have achieved considerable performance, several problems still exist. Firstly, it is challenging for supervised classification methods to train an accurate binary classifier due to the distribution mismatch between the expanded samples and the original samples. Secondly, it is difficult for unsupervised detection methods to design a discrimination loss that ensures effective background-target separation. Thirdly, it is demanding for both categories of methods to develop an efficient backbone to obtain robust intrinsic feature representations for HTD.
@@
Current backbones for HTD are mainly based on convolutional neural networks (CNNs) and Transformers. CNN-based methods [19, 20, 21, 27] excel at extracting local features but suffer from a limited receptive field, making it difficult to capture non-local correlations. In contrast, Transformer-based methods [23, 29, 30, 31] can capture global contextual dependencies based on the multihead self-attention mechanism. However, the computational complexity of Transformers is quadratic to the number of embedding tokens, leading to substantial computational overload and memory requirements as the sequence grows. Recently, structured state space models (SSMs) [32] have gained attention for modeling long sequences with near-linear complexity. Notably, Mamba [33] introduces a selective state space (S6) that makes SSM parameters input-aware, allowing for the selective propagation or forgetting of information based on the current token. Additionally, a hardware parallel scan algorithm in recurrent mode accelerates the training and inference process. Therefore, considering the high-dimensional sequence nature of hyperspectral data, applying Mamba as a backbone for HTD shows promise in efficiently capturing long-range dependencies.
@@
Inspired by the above insights, we propose an efficient HTD method with a pyramid SSM (HTD-Mamba). This method employs spectrally contrastive learning to recognize pixel-wise instances by maximizing the feature similarity between each pixel and its transformed view. Thus, detection can be performed by feeding pairs consisting of each detected pixel and the prior target spectrum into the well-trained model. Specifically, an efficient data augmentation technique is proposed to obtain sufficient training samples and leverage spatial information. This technique encodes all surrounding pixels within a patch into a new spectral view of the center pixel, using spectral similarity as weights. Additionally, to explore global band correlations, the input pixels are dynamically divided into continuous group-wise spectral embeddings using a one-dimensional (1D) CNN. The efficient architecture, Mamba, is first introduced to HTD to model long-range dependencies of the spectral sequence with linear complexity. Furthermore, to enhance robust feature representations, a pyramid SSM is developed as a backbone to extract multiresolution spectral-wise intrinsic features by capturing the global correlation of spectral sequences at different resolutions and fusing features from different levels. Therefore, HTD-Mamba effectively addresses the challenges of limited prior knowledge by constructing sufficient view pairs and mitigates spectral variation by extracting multiresolution discriminative intrinsic features using the pyramid SSM. Experimental results across multiple datasets demonstrate the effectiveness and superiority of HTD-Mamba compared to existing state-of-the-art methods.
@@
  2. We propose an efficient data augmentation technique to create sufficient view pairs and leverage spatial contextual information, which encodes all surrounding pixels within a patch into a transformed view of the center pixel.
@@ II-A1 SSM
The SSM is known as a linear time-invariant system that maps a 1D function or sequence $x(t)\in\mathbb{R}$ to its response $y(t)\in\mathbb{R}$. Mathematically, this process can be formulated using the following linear ordinary differential equations (ODEs):
@@
where $h(t)\in\mathbb{R}^{D}$ denotes the hidden state with size $D$, $h^{\prime}(t)$ denotes the time derivative of $h(t)$, ${\bf A}\in\mathbb{R}^{D\times D}$ is the state transition matrix derived from the high-order polynomial projection operator to retain historical information and capture long-range dependencies, and ${\bf B}\in\mathbb{R}^{D\times 1}$ and ${\bf C}\in\mathbb{R}^{1\times D}$ are the projection matrices controlling the input and output of the system, respectively.
Figure 1: The overall architecture of the proposed HTD-Mamba.
Figure 2: Flowchart of pyramid SSM.
@@ II-A2 Discretization
To integrate the continuous SSM into deep learning algorithms, the discretization process is necessary. Specifically, the ODEs (1) can be discretized using a zero-order hold rule as follows:
$\begin{split}h_{t}&=\bar{{\bf A}}h_{t-1}+\bar{{\bf B}}x_{t},\\ y_{t}&={\bf C}h_{t}\end{split}$ (2)
@@
$\begin{split}\bar{{\bf A}}&={\rm exp}({\bf\Delta}{\bf A}),\\ \bar{{\bf B}}&=({\bf\Delta}{\bf A})^{-1}({\rm exp}({\bf\Delta}{\bf A}-{\bf I}))\cdot{\bf\Delta}{\bf B}\end{split}$ (3)
@@ II-A3 Selective Scan Mechanism
Due to the time-invariant property, i.e., $\bf B$, $\bf C$, and ${\bf\Delta}$ are independent of the input sequence, the output can be calculated either through recurrance or by global convolution. However, this approach limits context awareness. To address this limitation, a selective scan mechanism is incorporated, where the projection matrices and time scale are derived from the input, ensuring awareness of the contextual information:
$\begin{split}S_{\bf B}({\bf z})&={\rm Linear}_{N}({\bf z}),\\ S_{\bf C}({\bf z})&={\rm Linear}_{N}({\bf z}),\\ S_{\bf\Delta}({\bf z})&={\rm Broadcast}_{D}({\rm Linear}_{1}({\bf z}))\end{split}$ (4)
where ${\rm Linear}_{q}$ represents the parameterized projection to dimension $q$, ${\rm Broadcast}_{q}$ denotes the broadcast operation to dimension $q$, and $N$ is the embedding size. Thus, the system becomes time-varying, and only recurrence can be adopted. Additionally, with the incorporation of parallel scanning and integration with the gated multilayer perceptron (MLP), the SSM evolves into the popular Mamba model, which excels at capturing long-range dependencies with near-linear time complexity and hardware acceleration.
@@ II-B Overall Architecture
To address the issues of limited training samples and spectral variation, we propose a self-supervised spectrally contrastive learning method based on Mamba with a pyramid SSM. This method aims to identify pixel-wise instances by maximizing the similarity and consistency between each pixel and its transformed view. Thus, targets of interest can be detected by pairing them with the prior target spectrum. Fig. 1 illustrates the overall architecture of the proposed method.
The training stage encompasses spatial-encoded spectral augmentation (SESA), multiresolution spectral feature extraction (MSFE), and spectral contrastive head (SCH). The SESA block initially creates a new spectral view of the center pixel by encoding contextual information within a patch, thereby generating sufficient view pairs for contrastive learning. Subsequently, the MSFE block divides the pixel into continuous group-wise spectral embeddings, treating them as a spectral sequence. To achieve discriminative intrinsic feature representation, we propose a pyramid SSM that captures multiresolution long-range dependencies within sequences. Finally, the paired extracted features pass through the SCH block to compute similarity, and contrastive loss is applied to ensure model discrimination by pulling positive pairs closer while pushing negative pairs farther apart.
In the testing stage, each detected pixel is paired with the prior target spectrum and then passes through the MSFE and SCH blocks to obtain the discriminative intrinsic feature vectors. The detection result of the $i$-th pixel is obtained by calculating the cosine similarity:
$\mu_{i}=c(f({\bf X}_{i}),f({\bf d}))$ (5)
where $c({\bf a},{\bf b})=({\bf a}^{T}{\bf b})/(\|{\bf a}\|_{2}\cdot\|{\bf b}\|_{2})$ denotes the cosine similarity function, $\|\cdot\|_{2}$ denotes the L2 norm, ${\bf X}\in\mathbb{R}^{n\times B}$ denotes the image matrix with $n$ pixels and $B$ bands, ${\bf d}\in\mathbb{R}^{B}$ denotes the target spectrum, and $f$ denotes the backbone consisting of MSFE and SCH. Finally, a nonlinear background suppression (NBS) function is applied to suppress the background:
@@ II-C Spatial-Encoded Spectral Augmentation
Given a random patch ${\bf P}\in\mathbb{R}^{p^{2}\times B}$, where the patch size is $p\times p$, and the center pixel is ${\bf y}\in\mathbb{R}^{B}$. By encoding all pixels within the patch using spectral similarity, the new spectral view $\bf x$ can be expressed as
${\bf x}={\bf P}^{T}{\bf W}$ (7)
where ${\bf W}\in\mathbb{R}^{p^{2}\times 1}$ is the weight matrix measuring the similarity between the center pixel and each pixel within the patch. The similarity weight corresponding to the $i$-th pixel within the patch is defined as
${\bf W}_{i}=\frac{{\rm exp}(c({\bf y},{\bf P}_{i}))}{\sum\limits_{j=1}^{p^{2}}{\rm exp}(c({\bf y},{\bf P}_{j}))}.$ (8)
After obtaining the new spectral view of the center pixel, pairs $({\bf x},{\bf y})$ from the same location comprise the positive samples while pairs from different locations comprise the negative samples. As a result, sufficient training samples can be obtained for contrastive learning.
@@ II-D Multiresolution Spectral Feature Extraction
Specifically, we use 1D convolution with a stride to extract local features and divide the spectrum into group representations:
@@
where ${\rm Conv}$ denotes the 1D convolution, $m$ denotes the kernel size, $s={\rm max}(1,\lfloor m/4\rfloor)$ denotes the stride, $N$ denotes the number of kernels, and $\sigma_{1}$ is the leaky rectified linear unit (Leaky ReLU)[34] function for nonlinear representation. After this process, paired spectral samples $({\bf x},{\bf y})$ can be converted to group-wise embeddings $({\bf S}_{x}\in\mathbb{R}^{L\times N},{\bf S}_{y}\in\mathbb{R}^{L\times N})$, where $L$ denotes the number of spectral groups, satisfying $L=\lfloor(B-m)/s\rfloor+1$. Thus, numerous spectral tokens can be obtained to extract discriminative intrinsic features.
@@
The flowchart of the proposed pyramid SSM is shown in Fig. 2. It comprises bottom-up downsampling, global spectral feature extraction, top-down upsampling and feature fusion, and the gated block. In each deep layer, the input sequence first undergoes a normalization operation to reduce the internal covariate shift and stabilize the training process. In this work, the root mean square layer normalization (RMSNorm)[35] is applied to improve the computational efficiency by eliminating the mean calculation required in the layer normalization:
@@
where ${\rm RMS}$ denotes the root mean square function, ${\bf S}_{i-1}^{j}$ denotes the $j$-th channel of the $i$-th layer input sequence, and ${\bf g}$ denotes the learnable weight. Then, the normalized sequence is projected into two representations for feature extraction and gated operation, each using a fully connected (FC) layer:
@@
In the bottom-up downsampling process, the spectral sequence is gradually downsampled to obtain multiresolution representations, which facilitates comprehensive analysis. Specifically, the adjacent hierarchical sampling is performed using convolution with stride:
@@
where ${\bf Z}_{1}^{0}={\bf Z}_{1}$, $k\in\{1,2,3\}$ denotes the $k$-th pyramid level. The kernel size is fixed at 3, the stride is 2 to perform 2$\times$ downsampling, and the token size is doubled to maintain information capacity.
@@
where $k\in\{0,1,2,3\}$, ${\rm DWConv}$ denotes the 1D depth-wise convolution, the kernel size is 3, the number of groups is $2^{k+1}N$, $\sigma_{2}$ denotes the sigmoid-weighted linear unit (SiLU)[36] function, and padding is used in the convolution to keep the size fixed.
In the top-down upsampling and feature fusion process, low-level features are recovered from the adjacent high level and then fused with the extracted global features from the same level. This process can be expressed as
@@
where $k\in\{3,2,1\}$, ${\hat{\bf Z}_{1}^{3}}={\bar{\bf Z}_{1}^{3}}$, ${\rm DConv}$ denotes the transposed 1D convolution with a kernel size of 3 and a stride of 2 to perform 2×2$\times$2 × upsampling. The token size is halved to maintain information capacity. The FC layer is used to perform flexible linear transformation in preparation for fusion.
@@
6:${\bar{\bf A}}$: $(P,L,N,D)$leftarrow ${\bf\Delta}\otimes{\rm Parameter}^{\bf A}$
7:${\bar{\bf B}}$: $(P,L,N,D)$leftarrow ${\bf\Delta}\otimes{\bf B}$
@@
10:for $k=1$ to 3 do
11:     $L_{k}=L_{k-1}//2$
12:     ${\bf Z}_{1}^{k}$: $(P,L_{k},2^{k+1}N)$leftarrow ${\rm Conv}({\bf Z}_{1}^{k-1})$
[…]
14:end for
15:/* Upsampling and feature fusion */
16:${\hat{\bf Z}_{1}^{3}}={\bar{\bf Z}_{1}^{3}}$
17:for $k=3$ to 1 do
18:     ${\hat{\bf Z}_{1}^{k-1}}$: $(P,L_{k-1},2^{k}N)$leftarrow ${\rm DConv}({\hat{\bf Z}_{1}^{k}}))$
19:     ${\hat{\bf Z}_{1}^{k-1}}$: $(P,L_{k-1},2^{k}N)$leftarrow ${\hat{\bf Z}_{1}^{k-1}}+{\rm Linear}_{2^{k}N}({\bar{\bf Z}_{1}^{k-1}})$
20:end for
21:/* Gated block */
22:${\bf Z}$: $(P,L,2N)$leftarrow ${\hat{\bf Z}_{1}^{0}}\odot\sigma_{2}({\bf Z}_{2})$
23:/* Residual connection */
24:${\bf S}_{i}$: $(P,L,N)$leftarrow ${\bf S}_{i-1}+{\rm Linear}_{N}({\bf Z})$
25:${\rm Return}:{\bf S}_{i}$
@@
An overview of the proposed pyramid SSM is summarized in Algorithm 1 and Algorithm 2, where $P$ denotes the batch size, $\otimes$ denotes the outer product, and ${\rm Parameter}^{\bf o}$ denotes the learnable parameter.